### v5

@@ 6 Supplementary Information: Super-additive Objective Functions
$c\left(\{X_{1},X_{2}\},\ldots\right)>c\left(X_{1},\ldots\right)+c\left(X_{2},\ldots\right).$ (22)
@@
[…]
$c\left(X_{1},\boldsymbol{X}^{\prime},\ldots\right)+c\left(X_{2},\boldsymbol{X}^{\prime\prime},\ldots\right)$ (23)
@@
7 Supplementary Information: Simple Example
@@
Note that all three optimal solutions place nodes $D$ and $G$ in the same cluster (cluster 2, shaded black in Fig. 1). $D$ is not only the node with the highest in-degree, but also connects the top part of the graph with the middle and, via node $G$, the bottom part of the graph. Thus it is feasible for $D$ and $G$ to be in the same cluster. Nodes $B$, $C$ and $E$ (shades of gray in Fig. 1) are all in individual clusters, corresponding to root ($B$, $C$) nodes or a node along a linear chain ($E$). As the objective function is based on computation cost, it is expected that root nodes are best assigned to their own clusters (Lemma 6.1). The key variation between the three least-cost cluster mappings is the mapping of node $A$ which can either be part of cluster 2 with $D$ and $G$, part of cluster 1 with node $F$, or in an independent cluster of its own. These solutions were discovered at iterations 35, 182, and 276 on branches 21, 27 and 104, respectively. With a cost of 54.0, this is much less than the naive heuristic solution cost of 85.8.
@@
Although the earlier solutions are quite different to the optimal solution, the solution at iteration 21 starts to follow the trend of the optimal solution with $D$, $G$ and $A$ in the same cluster, with other nodes mostly in their own clusters. This suggests that finding near-optimal or even the first optimal solution could potentially be very rapid. As might be expected, however, finding all optimal solutions could be expensive [17, 2]. In addition, near-optimal solutions could produce diverse cluster mappings, although they converge towards optimal mapping patterns fairly rapidly.
Declarations
  • Competing Interests: There are no competing interests.

### v6

@@ Appendix B Appendix: Super-additive Objective Functions
(24) $c\left(\{X_{1},X_{2}\},\ldots\right)>c\left(X_{1},\ldots\right)+c\left(X_{2},\ldots\right).$
@@
$\text{therefore, }c\left(\{X_{1},X_{2},\boldsymbol{X}^{\prime},\boldsymbol{X}^{\prime\prime}\},\ldots\right)>$
(25) $c\left(X_{1},\boldsymbol{X}^{\prime},\ldots\right)+c\left(X_{2},\boldsymbol{X}^{\prime\prime},\ldots\right)$
@@
Appendix C Appendix: Simple Example
@@
Note that all three optimal solutions place nodes $D$ and $G$ in the same cluster (cluster 2, shaded black in Fig. 6). $D$ is not only the node with the highest in-degree, but also connects the top part of the graph with the middle and, via node $G$, the bottom part of the graph. Thus it is feasible for $D$ and $G$ to be in the same cluster. Nodes $B$, $C$ and $E$ (shades of gray in Fig. 6) are all in individual clusters, corresponding to root ($B$, $C$) nodes or a node along a linear chain ($E$). As the objective function is based on computation cost, it is expected that root nodes are best assigned to their own clusters (Lemma B.1). The key variation between the three least-cost cluster mappings is the mapping of node $A$ which can either be part of cluster 2 with $D$ and $G$, part of cluster 1 with node $F$, or in an independent cluster of its own. These solutions were discovered at iterations 35, 182, and 276 on branches 21, 27 and 104, respectively. With a cost of 54.0, this is much less than the naive heuristic solution cost of 85.8.
@@
Although the earlier solutions are quite different to the optimal solution, the solution at iteration 21 starts to follow the trend of the optimal solution with $D$, $G$ and $A$ in the same cluster, with other nodes mostly in their own clusters. This suggests that finding near-optimal or even the first optimal solution could potentially be very rapid. As might be expected, however, finding all optimal solutions could be expensive (Likhachev2005; Belfodil2019). In addition, near-optimal solutions could produce diverse cluster mappings, although they converge towards optimal mapping patterns fairly rapidly.
Appendix D Reproducibility Checklist for JAIR
Select the answers that apply to your research – one per item.
All articles:
  1. (1)
    All claims investigated in this work are clearly stated. [yes]
  2. (2)
    Clear explanations are given how the work reported substantiates the claims. [yes]
  3. (3)
    Limitations or technical assumptions are stated clearly and explicitly. [yes]
  4. (4)
    Conceptual outlines and/or pseudo-code descriptions of the AI methods introduced in this work are provided, and important implementation details are discussed. [yes]
  5. (5)
    Motivation is provided for all design choices, including algorithms, implementation choices, parameters, data sets and experimental protocols beyond metrics. [yes]
Articles containing theoretical contributions:
Does this paper make theoretical contributions? [yes]
If yes, please complete the list below.
  1. (1)
    All assumptions and restrictions are stated clearly and formally. [yes]
  2. (2)
    All novel claims are stated formally (e.g., in theorem statements). [yes]
  3. (3)
    Proofs of all non-trivial claims are provided in sufficient detail to permit verification by readers with a reasonable degree of expertise (e.g., that expected from a PhD candidate in the same area of AI). [yes]
  4. (4)
    Complex formalism, such as definitions or proofs, is motivated and explained clearly. [yes]
  5. (5)
    The use of mathematical notation and formalism serves the purpose of enhancing clarity and precision; gratuitous use of mathematical formalism (i.e., use that does not enhance clarity or precision) is avoided. [yes]
  6. (6)
    Appropriate citations are given for all non-trivial theoretical tools and techniques. [yes]
Articles reporting on computational experiments:
Does this paper include computational experiments? [yes]
If yes, please complete the list below.
  1. (1)
    All source code required for conducting experiments is included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for source code readability and documentation as well as for long-term accessibility. [yes]
  2. (2)
    The source code comes with a license that allows free usage for reproducibility purposes. [yes]
  3. (3)
    The source code comes with a license that allows free usage for research purposes in general. [yes]
  4. (4)
    Raw, unaggregated data from all experiments is included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for long-term accessibility. [yes]
  5. (5)
    The unaggregated data comes with a license that allows free usage for reproducibility purposes. [yes]
  6. (6)
    The unaggregated data comes with a license that allows free usage for research purposes in general. [yes]
  7. (7)
    If an algorithm depends on randomness, then the method used for generating random numbers and for setting seeds is described in a way sufficient to allow replication of results. [NA]
  8. (8)
    The execution environment for experiments, the computing infrastructure (hardware and software) used for running them, is described, including GPU/CPU makes and models; amount of memory (cache and RAM); make and version of operating system; names and versions of relevant software libraries and frameworks. [yes]
  9. (9)
    The evaluation metrics used in experiments are clearly explained and their choice is explicitly motivated. [yes]
  10. (10)
    The number of algorithm runs used to compute each result is reported. [yes]
  11. (11)
    Reported results have not been “cherry-picked” by silently ignoring unsuccessful or unsatisfactory experiments. [yes]
  12. (12)
    Analysis of results goes beyond single-dimensional summaries of performance (e.g., average, median) to include measures of variation, confidence, or other distributional information. [yes]
  13. (13)
    All (hyper-) parameter settings for the algorithms/methods used in experiments have been reported, along with the rationale or method for determining them. [yes]
  14. (14)
    The number and range of (hyper-) parameter settings explored prior to conducting final experiments have been indicated, along with the effort spent on (hyper-) parameter optimisation. [yes/]
  15. (15)
    Appropriately chosen statistical hypothesis tests are used to establish statistical significance in the presence of noise effects. [NA]
Articles using data sets:
Does this work rely on one or more data sets (possibly obtained from a benchmark generator or similar software artifact)? [no]
If yes, please complete the list below.
  1. (1)
    All newly introduced data sets are included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for long-term accessibility with a license that allows free usage for research purposes. [yes/partially/no/NA]
  2. (2)
    The newly introduced data set comes with a license that allows free usage for reproducibility purposes. [yes/partially/no]
  3. (3)
    The newly introduced data set comes with a license that allows free usage for research purposes in general. [yes/partially/no]
  4. (4)
    All data sets drawn from the literature or other public sources (potentially including authors’ own previously published work) are accompanied by appropriate citations. [yes/no/NA]
  5. (5)
    All data sets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. [yes/partially/no/NA]
  6. (6)
    All new data sets and data sets that are not publicly available are described in detail, including relevant statistics, the data collection process and annotation process if relevant. [yes/partially/no/NA]
  7. (7)
    All methods used for preprocessing, augmenting, batching or splitting data sets (e.g., in the context of hold-out or cross-validation) are described in detail. [yes/partially/no/NA]
Explanations on any of the answers above (optional):
The cost function used in the case study only depends on the parent-child relationships in the directed acyclic graph, and the number of states of each discrete node. Hence, the only data for reproducibility is based on that to reproduce the case study examples.
The results are independent of the machine running the algorithm. Nevertheless, we used a Windows 11 PC running 14900HX CPU with 32GB of RAM and using the Armadillo library version 12.8.4.